OpenStudy (anonymous):

x/4 + 1= -5 -2 16 4 -24 I thought it would be -16 but that's not one of the choices

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x}{4}+1 = -5\] \[\Large \frac{x}{4} = -5-1\] \[\Large \frac{x}{4} = -6\] See how to get the answer now?

OpenStudy (bibby):

you want to isolate the x first. subtract one and multiply both sides by 4. -24.

jimthompson5910 (jim_thompson5910):

you mistakenly did -5-1 = -4, but that's not correct since it should be -5-1 = -6

OpenStudy (anonymous):

what?

jimthompson5910 (jim_thompson5910):

you first subtract 1 from both sides that will give you \[\Large \frac{x}{4} = -6\] and NOT \[\Large \frac{x}{4} = -4\]